OpenStudy (anonymous):

Derivative of cos^2x The chain rule gets a different answer than the power reducing formula

zepdrix (zepdrix):

`The chain rule gets a different answer than the power reducing formula` What? 0_o You must apply both rules, power rule first.

zepdrix (zepdrix):

Oh you applying Half-Angle Formula to cosine before differentiating? Maybe I misunderstood :)

zepdrix (zepdrix):

blah I'll wait ^^

OpenStudy (anonymous):

If the chain rule is used to find\[\frac{ d }{ dx } \cos^2x = 2\cos(x)(\cos)'=> -2cosx(sinx)\]

OpenStudy (anonymous):

If the power reducing formula is used \[\cos^2x = \frac{ 1 + \cos 2x }{ 2 }\] \[\frac{ d }{ dx } \frac{ 1 + \cos2x }{ 2 } = -sinx \] Did I make a mistake somewhere?

zepdrix (zepdrix):

\[\large\rm \frac{d}{dx}\frac{1}{2}+\frac{1}{2}\cos(2x)\quad=0+\frac{1}{2}(-\sin(2x))(2x)'\]

zepdrix (zepdrix):

\[\large\rm =-\sin(2x)\]Ya? :)

zepdrix (zepdrix):

cosine turns into minus sine, (The inner function shouldn't change at this point, the angle is still 2x) then chain rule

zepdrix (zepdrix):

So it's this,\[\large\rm \frac{d}{dx}\cos(2x)\quad\approx-\sin(2x)\]but with chain rule,\[\large\rm \frac{d}{dx}\cos(2x)\quad=-\sin(2x)(2x)'\]\[\large\rm \frac{d}{dx}\cos(2x)\quad=-2\sin(2x)\]Confusing? :o

OpenStudy (anonymous):

Using the power reducing formula y' = -sin(2x) But using the chain rule for cos^2x it is −2cosx(sinx)

zepdrix (zepdrix):

Good observation! :) Recall your Sine Double-Angle Formula:\[\large\rm \sin(2x)=2sinxcosx\]

OpenStudy (anonymous):

Thank you very much!

zepdrix (zepdrix):

np